OpenStudy (anonymous):

Pyruvate has the chemical formula C3H3O3. Can someone please explain to me how to draw pyruvate with a pH of 7.4?? Thank you :)

OpenStudy (aaronq):

the 2nd link is for methyl pyruvate, not pyruvate. this is pyruvic acid:  using the henderson-hasselbalch equation, we can find if the acidic H has dissociated or not. \(pH=pKa+log\dfrac{[A^-]}{[HA]}\) \(7.4=2.5+log\dfrac{[A^-]}{[HA]}\) \(\dfrac{[A^-]}{[HA]}=\dfrac{79432.823}{1}\) we can see that its present mostly in the deprotonated form, so, this is what it looks like at pH 7.4 

OpenStudy (anonymous):

Thank you @aaronq